The maintenance cost for a generator have been recorded over its five year life. Calculate EUAC at 8 percent per year.

Year

1

2

3

4

5

Cost

$1100

1500

1300

1900

2200
Answer #1

Given

Rate of interest=i=8%

We first estimate the PV of future costs.

PV=-[1100*(F/P,8%,1)+1500*(F/P,8%,2)+1300*(F/P,8%,3)+1900*(F/P,8%,4)+2200*(F/P,8%,5)]

Let us calculate the interest factors

(F/P,8%,1)=1/(1+8%)^1=0.925926

(F/P,8%,2)=1/(1+8%)^2=0.857339

(F/P,8%,1)=1/(1+8%)^3=0.793832

(F/P,8%,1)=1/(1+8%)^4=0.735030

(F/P,8%,1)=1/(1+8%)^1=0.680583

PV=-[1100*0.925926+1500*0.857339+1300*0.793832+1900*0.735030+2200*0.680583]=-$6230.3483

EUAC=PV*(A/P,8%,5)

Let us calculate the interest factor.

EUAC=PV*(A/P,8%,5)=-6230.3483*0.250456 =-$1560.43